# Stein Variational Evolution Strategies

### v1

@@ 5 Experiments
We compare SV-CMA-ES with the two existing gradient-free SVGD algorithms from literature: GF-SVGD (Han and Liu, 2018) and the MC gradient SVGD (Liu et al., 2017) that we refer to as SV-OpenAI-ES following the naming convention of the ES community. Furthermore, we compare against vanilla SVGD as a gradient-based method. All strategies have been implemented based on the evosax (Lange, 2023) library. To guarantee a fair comparison, we keep the number of function evaluations equal across all methods. In other words, if SV-CMA-ES is evaluated for 4 particles, each represented by a subpopulation of size 16, we evaluate GF-SVGD and SVGD with 64 particles. For each kernel-based method, we use the standard Gaussian RBF kernel. For GF-SVGD, we follow the setup of Han and Liu (2018) and use the same kernel function for the SVGD kernel $k$ and the smoothing kernel $k_{\rho}$, as well as an isotropic Gaussian prior ${\cal N}(0,\sigma^{2}{\rm\bf I})$. The optimal scale $\sigma^{2}$ of the prior is found via a hyperparameter grid search. For each method, we implement the annealed version of SVGD (cf. Eq. (3)) using a logarithmic schedule $\gamma(t)=\max(\log(T/t),1)$. Unless specified differently, this choice is followed in all experiments. For all optimizer-based methods, we use Adam (Kingma and Ba, 2015). For a fair comparison, we search the best hyperparameters for all algorithms separately. The full details of our experimental setup can be found in the Appendix B and in our codebase. [^1] We will provide a link to our code here upon conference publication.
@@ 5.1 Sampling from Synthetic Densities
Setting
We first evaluate or method on multiple synthetic densities to illustrate the quality of the generated samples. The closed form pdf for every problem is listed in the Appendix. We use the total population size of 400 across all methods. For the ES, this number is split across 100 particles, while for SVGD and GF-SVGD the total population size is equal to the number of particles. We additionally evaluate the scaling of all methods. For this analysis, we report the final performances after 1000 iterations across different numbers of particles, as well as subpopulation sizes for the ES based methods.
Gaussian Mix.
@@
(b) SVGD
(c) GF- SVGD
(d) SV- OpenAI-ES
(e) SV-CMA- ES
Figure 2: Samples obtained by various methods after 1000 iterations. Gradient-based SVGD captures all target densities effectively, while SV-CMA-ES produces the highest quality samples among gradient-free methods. GF-SVGD struggles with complex distributions, and SV-OpenAI-ES tends to converge slowly due to taking small steps in flat regions of the target.
Results
The Figures 2 and 3 display the qualitative and quantitative sample properties. As expected, gradient-based SVGD generates high quality samples for all problems. We find that among the gradient-free methods, SV-CMA-ES performs the best across all problems. While GF-SVGD generates high quality samples for the Gaussian mixture, the variance of the generated samples on the double banana density is too high, and the samples for the motion planning problem are of general bad quality. Similarly, SV-OpenAI-ES performs well on the motion planning problem, but on the others it underestimates the variance (Gaussian mixture) or converges slowly (double banana). We hypothesize that GF-SVGD fails on the motion planning problem because choosing a good prior for this problem is difficult. These results highlight the fast convergence properties of SV-CMA-ES, as it employs the automatic step-size adaptation of CMA-ES. In a further analysis, we investigate the performances for different particle numbers and population sizes on the same problems as above. The results of these experiments are displayed in Figure 5. In addition to the Maximum Mean Discrepancy (Gretton et al., 2012, MMD), we report the error when estimating the first two central moments of the target distribution from the generated samples. We observe that there is a clear trend that SV-CMA-ES performs better than GF-SVGD and SV-OpenAI-ES with an increasing number of particles. Furthermore, we note in Fig. 5 (d) that SV-CMA-ES requires fewer samples than SV-OpenAI-ES to estimate good steps.
@@ Results
(d) Aggregated MMD
Figure 3: (a)-(c): MMD convergence results on the synthetic densities depicted in Fig. 2. (d) Mean log10 MMD across all three sampling tasks. All results are averaged across 10 independent runs (±$\pm 1.96$ standard error). We observe that SV-CMA-ES produces the best samples out of the two gradient-free baselines and is competitive with gradient-based SVGD (blue line) on the presented tasks.
@@ 5.2 Bayesian Logistic Regression
Next, we evaluate our method on Bayesian logistic regression for binary classification. We follow the setup described by Langosco di Langosco et al. (2021), which uses a hierarchical prior $p(\theta)$ on the parameters $\theta=[\alpha,\beta]$, where $\beta\sim{\cal N}(0,\alpha^{-1})$ and $\alpha\sim\Gamma(a_{0},b_{0})$. Given this prior, the task is to approximate samples from the posterior $p(\theta\mid D)=p(D\mid\theta)p(\theta)$ with:
@@
We consider the binary Covtype and the Glass datasets (Asuncion et al., 2003), as suggested in the original SVGD and GF-SVGD works (Liu and Wang, 2016; Han and Liu, 2018). We partition the data into 70% for training, 10% for validation and 20% for testing and use a batch size of 128. For all experiments, we use a total population of 256, which is split across 8 particles for the ES.
@@ Results
Our results demonstrate SV-CMA-ES outperforms the remaining gradient-free optimizers. On both datasets, our methods converges faster than SV-OpenAI-ES and GF-SVGD. Furthermore, its performance is considerably better than GF-SVGD for the Covtype dataset. We want to highlight that SV-CMA-ES achieves these results despite using only a fraction of the number of particles compared to the single population baselines, because each particle requires a subpopulation size of 32. Running SVGD with fewer particles typically reduces the optimizers chances to collect good samples. Nevertheless, SV-CMA-ES performs similar to gradient-based SVGD.
@@
(b) Glass
Figure 4: Results of Bayesian logistic regression on the Covtype and Glass datasets. We report mean (±$\pm 1.96$ standard error intervals across 10 independent runs. SV-CMA-ES converges the faster than other gradient-free methods, and performs similarly well as gradient-based SVGD (dashed line).
(a) MMD
(b) $\operatorname*{\mathbb{E}}[x]$
(c) ${\mathbb{V}}[x]$
(d) Subpopulation Scaling
Figure 5: Scaling analysis. Depicted are the final performances for different total population sizes. (a) MMD vs. sample size after 1000 iterations. (b)-(c) MSE vs sample size when estimating the first two central moments of the target distribution. For ES we use the same subpop. size as in Fig. 3. (d) Subpopulation size scaling for ES-based SVGD. The results are averaged across 10 independent of all synthetic sampling tasks from Fig. 3. SV-CMA-ES performs the best out of all gradient-free methods (solid lines) across different particle numbers.
@@ 5.3 Reinforcement Learning
We further assess the performances of the gradient-based SVGD methods on six classic reinforcement learning (RL) problems. The goal of each RL task is to maximize the expected episodic return $J(\theta)$, where each particle $\theta$ now parameterizes a multi-layer perceptron (MLP) controller. The corresponding SVGD objective is to sample from the following posterior:
$p(\theta)\propto\exp(-J(\theta)),\quad J(\theta)=\operatorname*{\mathbb{E}}_{(s_{t},a_{t})\sim\pi_{\theta}}\bigl{[}\sum_{t=1}^{T}R(s_{t},a_{t})\bigr{]}$
where $(s_{t},a_{t})\sim\pi_{\theta}$ represent a trajectory sampled from the distribution that is induced by the policy parametrized by $\theta$. For each problem, we train a 2-hidden layer MLP with 16 units per layer. We use a total population size of 64 which we split into 4 subpopulations for the ES-based methods and estimate the expected return across 16 rollouts with different seeds. To make the results comparable to other works on ES for RL, we follow the approach of Lee et al. (2023) and extend the optimization by a phase that attempts to find exact optima. We realize this by fading out the repulsive term via the schedule $\gamma(t)=\log(T/t)$.
Results
Our results showcase strong performances of SV-CMA-ES. In comparison to other gradient-free versions of SVGD, it is the only method that is capable of obtaining high rewards on every problem that we test on. In particular, we observe that SV-CMA-ES is the only method that solves the MountainCar problem consistently, while it converges the fastest on Pendulum. Both of these environments feature a local optimum at which agents remain idle to avoid negative control costs (Eberhard et al., 2023). On the same problems, GF-SVGD converges to exactly these optima in some runs. These results illustrate that SV-CMA-ES improves over GF-SVGD by sampling stochastic SVGD steps, which leads to a better exploration of the domain. Interestingly, our results further show that SV-OpenAI-ES may deliver good samples in some runs, but the high standard error on several problems underline that it is sensitive to initialization and thus fails to perform well for some seeds. These findings further confirm that SV-CMA-ES is a strong gradient-free variant of SVGD, capable of both sampling from densities and optimizing blackbox objectives. We extend our analysis of SV-CMA-ES in the Appendix where we compare its performance for blackbox optimization to vanilla CMA-ES and OpenAI-ES, and conduct additional evaluations on sparse reward environments. These experiments reveal that SV-CMA-ES consistently outperforms competing algorithms, underscoring its superior capability in environments where effective exploration is essential.
@@ 6 Conclusion
We proposed a new gradient-free algorithm for Stein variational inference. The resulting method, SV-CMA-ES, combines elements from evolutionary computation and SVGD to sample from unnormalized densities without gradients. On several problems with different characteristics, we demonstrated that SV-CMA-ES outperforms prior gradient-free SVGD versions consistently. We could thus confirm our hypothesis that the incorporation of the CMA-ES update enables faster convergence than vanilla MC gradients, and better overall performance than GF-SVGD across multiple problems.
@@
For stable convergence, we selected a fixed kernel bandwidth via grid search, while prior work used the median heuristic. Selecting the kernel bandwidth via grid search is costly and thus constitutes a disadvantage. Furthermore, our approach can be computationally expensive due to the singular value decomposition required for each covariance matrix. The latter leads to a runtime complexity in $\mathcal{O}(nd^{3})$ in $d$ dimensions with $n$ particles. Future work could address this by exploring diagonal covariance matrices, which are commonly used to speed up CMA-ES (Ros and Hansen, 2008).
@@
In our experiments, we used the standard RBF kernel as it is convention in many prior works. One potential extension is to use SV-CMA-ES with matrix-valued kernels (Wang et al., 2019a) to explicitly consider the covariance matrix of the CMA-ES update. Our paper proposes a general framework of approximating SVGD steps using ES. The SV-CMA-ES algorithm that we introduced is one instance of this. We see it as an exciting avenue for future research to investigate other ES in this context, for instance PGPE (Sehnke et al., 2010) or NRES (Li et al., 2023).
Acknowledgments
This research was funded by the Amazon Development Center Germany GmbH.
Supplementary Materials
Appendix A Vanilla CMA-ES Algorithm
The listing in Algorithm 2 displays the update step of vanilla CMA-ES at iteration $t$. Parallel CMA-ES, on which we base our method, uses the same update, but carries it out $n$ times in parallel, once for each search distribution. Following the main paper, we denote the size of the CMA-ES population by $m\in{\mathbb{N}}^{+}$. For more details on how the weights $w_{i\dots m}$ and other hyperparameters can be chosen, we refer to Hansen (2016). In our experiments, we tune the elite population size $\lambda$, the initial value for $\sigma$ using a grid search (cf. Appendix B) and use the defaults from the evosax codebase (Lange, 2023) for the remaining parameters.
Algorithm 2 The Vanilla CMA-ES Update. Adapted from Hansen (2016).
Input: Generation index $t$, search distribution parameters $\mathbf{x},\sigma,\mathbf{C}$; elite pop. size $\lambda$; learning rates $\alpha_{\mathbf{x}},\alpha_{\sigma},\alpha_{1},\alpha_{\lambda},\alpha_{c}$; damping hyperparam. $d_{\sigma}$
1:Sample and evaluate new population of search points, for $i=1,\dots,m$
$\mathbf{y}_{i}=(\boldsymbol{\xi}_{i}-\mathbf{x})/\sigma,\quad\text{where }\boldsymbol{\xi}_{i}\sim{\cal N}(\mathbf{x},\sigma^{2}\mathbf{C})$ (15)
$f_{i}=f(\boldsymbol{\xi}_{i})$ (16)

### v2

@@ 5 Experiments
We compare SV-CMA-ES against the two existing approaches for zero-order SVGD from the literature: GF-SVGD as state-of-the-art method for surrogate-based inference, and the MC gradient SVGD as state-of-the-art gradient approximation method. We refer to the latter as SV-OpenAI-ES throughout the remainder of the paper following the naming convention of the ES community. Furthermore, we compare against gradient-based SVGD, which we denote as $\nabla$-SVGD in the following. All strategies have been implemented based on the evosax library (Lange, 2023).
To guarantee a fair comparison, we keep the number of function evaluations equal for all methods. In other words, if the ES-based methods are evaluated for 4 particles, each sampling subpopulations of size 16, we evaluate GF-SVGD and $\nabla$-SVGD with 64 particles. For each kernel-based method, we use the standard RBF kernel. For GF-SVGD, we follow the setup of Han and Liu (2018) and use the same kernel function for the SVGD kernel $k$ and the surrogate kernel $k_{\rho}$, as well as an isotropic Gaussian prior ${\cal N}(0,\sigma^{2}{\rm\bf I})$. The optimal scale $\sigma^{2}$ of the prior is found via a hyperparameter grid search. For each method, we implement the annealed version of $\nabla$-SVGD (cf. Eq. (3)) using a logarithmic schedule $\gamma(t)=\max(\log(T/t),1)$. Unless specified differently, this choice is followed in all experiments. For all methods that require an internal optimizer, we use Adam (Kingma and Ba, 2015). We carefully search for the best hyperparameters for each algorithm separately, to guarantee a fair comparison. The full details of our experimental setup can be found in the Appendix B. Moreover, we refer to Appendix C for additional results including ablation studies and empirical runtime and convergence analyses.
@@ 5.1 Sampling from Synthetic Densities
Gaussian Mixture
@@
(b) $\nabla$-SVGD
(c) SV-CMA-ES (ours)
(d) GF-SVGD
(e) SV- OpenAI-ES
Figure 2: Samples obtained by various methods. Gradient-based SVGD (b) captures all target densities effectively, while SV-CMA-ES produces the highest quality samples among gradient-free methods. GF-SVGD struggles on more complex targets, and SV-OpenAI-ES tends to converge slowly due to taking small steps in flat regions of the target.
@@
(d) MMD w.r.t. $\nabla$-SVGD
Figure 3: (a)-(c): MMD w.r.t. ground truth samples on the synthetic densities depicted in Figure 2. (d): Mean log10 MMD across all three sampling tasks w.r.t. the samples obtained by gradient-based SVGD. All results are averaged across 10 independent runs (±$\pm 1.96$ standard error). SV-CMA-ES approximates the ground truth samples and results by gradient-based SVGD (blue line) the best out of all gradient-free methods.
Setting
We first evaluate our method on multiple synthetic densities to illustrate the quality of the generated samples. The closed form pdf for every problem is listed in Appendix B. We use a total population size of 400 for all methods, which is split across 100 particles for the ES-based algorithms. In other words, each particle samples an ES population of 4. Following common practice in the literature, we quantify sampling performance by evaluating the Maximum Mean Discrepancy (Gretton et al., 2012, MMD) of the particles with respect to ground truth samples. We additionally evaluate the scaling to higher particle numbers in Sec. 5.4.
Results
Figures 2 and 3 display the qualitative and quantitative sampling results. As expected, $\nabla$-SVGD generates high-quality samples for all problems. We find that among the gradient-free methods, SV-CMA-ES performs the best across all problems. While GF-SVGD generates high-quality samples for the Gaussian mixture, the variance of the generated samples on the double banana density is too high, and the samples for the motion planning problem are of poor quality. Concurrently, SV-OpenAI-ES performs well on the motion planning problem, but on the others it underestimates the variance (Gaussian mixture) or converges slowly (double banana, also see Fig. 13). These results highlight the fast convergence properties of our method, as it employs the automatic step-size adaptation of CMA-ES. Additionally, Figure 3 (d) displays the MMD w.r.t. the samples that were obtained by $\nabla$-SVGD, aggregated across all sampling tasks. Our results demonstrate that SV-CMA-ES can indeed quickly converge to a set of samples that approximates the outcomes of $\nabla$-SVGD well, and better than the two other gradient-free baselines. We further illustrate these results in Section C.5 where we display the sample sets for all sampling tasks. Moreover, we illustrate the benefit of using the presented algorithm compared to other CMA-ES-based methods in Fig. 1 – since prior CMA-ES methods only maximize likelihood, the diversity of samples is low.
@@ 5.2 Bayesian Logistic Regression
Next, we evaluate our method on Bayesian logistic regression for binary classification. We follow the setup of Langosco et al. (2021), which uses a hierarchical prior $p(\theta)$ on the parameters $\theta=[\alpha,\beta]$, where $\beta\sim{\cal N}(0,\alpha^{-1})$ and $\alpha\sim\Gamma(a_{0},b_{0})$. Given data $D$, the task is to approximate samples from the posterior
$p(\theta\mid D)=p(D\mid\theta)p(\theta)\quad\text{ with:}$:
@@
We consider the binary Covtype, Spambase, and the German credit datasets from the UCI Machine Learning Repository (Kelly et al., 2023), as suggested in prior work (Liu and Wang, 2016; Arenz et al., 2020; Futami et al., 2018). For all experiments, we use a total population of 256, which is split across 8 particles for the ES-based methods.
@@ Results
For each dataset, we report the accuracy and negative log-likelihood (NLL) across the entire particle set, and report the mean performance across 10 runs. Our results demonstrate that SV-CMA-ES outperforms the remaining gradient-free algorithms. On both datasets, our method is the fastest converging among the gradient-free methods. Furthermore, its final performance is considerably better than GF-SVGD on all datasets. While the performance of $\nabla$-SVGD is slightly better on the Covtype dataset, SV-CMA-ES is on par with it for the Spam dataset. Additionally, on the credit data, we find that ES-based methods are both more accurate and exhibit greater stability than the gradient-based SVGD, which underlines the potential of zero-order methods in this context.
@@
(b) Spam
(c) Credit
Figure 4: Results of Bayesian logistic regression. We report mean (±$\pm 1.96$ standard error) across 10 independent runs. SV-CMA-ES converges the faster than other gradient-free methods, and achieves similar performance levels at convergence as gradient-based SVGD (dashed line).
@@ 5.3 Reinforcement Learning
We further assess the performances of the gradient-free SVGD methods on six classic reinforcement learning (RL) problems. The goal of each RL task is to maximize the expected episodic return $J(\theta)$, where each particle $\theta$ now parametrizes a multi-layer perceptron (MLP). The corresponding inference objective is to sample policy parameters $\theta$ from the following Boltzmann distribution:
$p(\theta)\propto\exp(J(\theta)),\quad J(\theta)=\operatorname*{\mathbb{E}}_{(s_{t},a_{t})\sim\pi_{\theta}}\bigl{[}\sum_{t=1}^{T}r(s_{t},a_{t})\bigr{]}$
where $(s_{t},a_{t})\sim\pi_{\theta}$ represent a trajectory sampled from the distribution that is induced by the policy parametrized by $\theta$. For each problem, we train a 2-hidden layer MLP with 16 units per layer, which implies high-dimensional optimization problems as each MLP has several hundred parameters. The specific numbers vary across the benchmarks and are listed in Table 2 in the Appendix. We use a total population size of 64 which we split into 4 subpopulations for the ES-based methods and estimate the expected return across 16 rollouts with different seeds. To make the results comparable to other works on ES for RL, we follow the approach of Lee et al. (2023) and extend the optimization by a phase that attempts to find exact optima. We realize this by fading out the repulsive term via the schedule $\gamma(t)=\log(T/t)$.
@@
Results
We display the aggregated results across all RL tasks in Fig. 1, and the individual task performances in Fig. 5. Our results showcase a strong performance of SV-CMA-ES. In comparison to other gradient-free versions of SVGD, it is the only method that generates high scoring solutions for all problems. In particular, we observe that SV-CMA-ES is the only method that solves the MountainCar problem consistently, while it is the fastest to converge on Pendulum. Both of these environments feature a local optimum at which agents remain idle to avoid control costs (Eberhard et al., 2023). It is on these problems that GF-SVGD converges to such optima in certain runs, which we further illustrate in Fig. 8 in the Appendix. These results illustrate that SV-CMA-ES improves over GF-SVGD by sampling stochastic ES steps, which leads to a higher exploration of the domain. Interestingly, our results further show that SV-OpenAI-ES may deliver good samples in some runs, but the high standard error on several problems underline its sensitivity to initialization. These findings confirm that SV-CMA-ES is a strong gradient-free SVGD scheme, capable of sampling from densities and optimizing blackbox objectives. Further, we would like to note that our final performances are comparable to those reported in prior, gradient-based, work (Jesson et al., 2024), which again underlines the potential of our method.
Further, we analyze the benefits of the kernel term by comparing our method to uncoordinated parallel runs of CMA-ES. Overall, we observe a clear performance improvement when using the kernel term. In particular, in the more challenging Hopper and Walker tasks, the benefits of using SV-CMA-ES over parallel CMA-ES are large. We extend our analysis of SV-CMA-ES in Appendix C.1 where we compare it to vanilla CMA-ES and OpenAI-ES, and conduct additional experiments on sparse reward environments. This analysis reveals that SV-CMA-ES consistently outperforms competing ES, underscoring its superior performance in environments where effective exploration is essential.
5.4 Ablation Studies
Choice of Population Size
In the experiments above, we investigate the performance for fixed particle numbers and population sizes. To gain further insights into the scalability of our method, we conduct an additional analysis on the same sampling problems as in Section 5.1 using varying particle numbers. The results of these experiments are displayed in Figure 6. In addition to the MMD after 1 00010001\,0001 000 iterations, we report the error when estimating the first two central moments of the target distribution from the generated samples. We observe the clear trend that SV-CMA-ES performs better than GF-SVGD and SV-OpenAI-ES with increasing particle numbers. Furthermore, we observe in Fig. 6 (d) that SV-CMA-ES requires fewer samples than SV-OpenAI-ES to estimate good steps.
Choice of Annealing Schedule
In the experiments above, we used annealed SVGD. This decision was made due to its widespread use in the community and many desirable properties. However, to assess the quality of our method, it is important to consider the sensitivity to the choice of annealing schedule. In Table 1, we show the key performance metrics from 10 seeds for SV-CMA-ES with and without annealing. As we see, performance in all cases is strong, with little difference between the two conditions.
(a) MMD w.r.t. ground truth
(b) $\operatorname*{\mathbb{E}}[x]$
(c) ${\mathbb{V}}[x]$
(d) Subpopulation Scaling
Figure 6: Scaling analysis. Depicted are the final performances for different total population sizes. (a): MMD vs. sample size after 1000 iterations. (b)-(c): MSE vs. sample size when estimating the first two central moments of the ground truth distribution. For ES, we use the same subpop. size per particle as in Figure 3. (d): Subpopulation size scaling for ES-based SVGD. The results are averaged across 10 independent runs of all synthetic sampling tasks from Figure 3. SV-CMA-ES performs the best out of all gradient-free methods (solid lines) across different particle numbers.
Task Annealing No Annealing
GMM -3.03 (0.30) -2.92 (0.26)
Double Banana -2.59 (0.16) -2.83 (0.21)
Motion Planning -2.40 (0.10) -2.44 (0.13)
Covertype NLL 0.59 (0.01) 0.59 (0.01)
MountainCar RL 93.68 (0.10) 93.68 (0.08)
Hopper RL 1781.31 (132.64) 1788.13 (79.22)
Table 1: Kernel annealing ablation. This table shows SV-CMA-ES performances across 10 seeds with 1.96 standard error in parentheses. All runs use the identical setup, aside from the kernel annealing. No annealing means that we use a constant $\gamma(t)=1$.
@@ 6 Conclusion
We proposed a new gradient-free algorithm that combines elements from evolution strategies and SVGD. The resulting method, SV-CMA-ES, achieves high computational efficiency by replacing the score term in the SVGD update with an ES step. On several problems with different characteristics, we demonstrated that SV-CMA-ES outperforms prior gradient-free SVGD-based algorithms consistently. We could thus confirm our hypothesis that the incorporation of the CMA-ES update enables faster convergence than SVGD with MC gradients, and better overall performance than GF-SVGD across multiple problems.
@@
For stable convergence, we selected a fixed kernel bandwidth via grid search, while prior work used the median heuristic. Selecting the kernel bandwidth via grid search is costly and thus constitutes a disadvantage. Furthermore, our approach can be computationally expensive due to the decomposition required for each covariance matrix, leading to a runtime complexity in $\mathcal{O}(\varrho^{2}d+\varrho d^{3})$ in $d$ dimensions with $\varrho$ particles. In contrast, SV-OpenAI-ES and GF-SVGD achieve a complexity in $\mathcal{O}(\varrho^{2}d)$. Future work could address this by exploring diagonal covariance matrices, which are commonly used to speed up CMA-ES (Ros and Hansen, 2008). Additionally, we would like to stress that the most time-consuming part of ES is often the fitness evaluation. We illustrate this aspect in Section C.3 where we present additional plots including empirical runtimes. This analysis shows that the wallclock time that SV-CMA-ES requires to produce high-quality solutions is competitive with the baselines.
@@
In our experiments, we used the standard RBF kernel, following the convention of many prior works. Recent work suggested adjusting the size of the considered neighborhood adaptively in the context of particle swarm optimization (Zhang et al., 2024). One potential extension is to integrate this idea into our approach to improve particle repulsion. Moreover, we see, for instance in Figure 3, that our method has higher variance compared to other methods. Future work could investigate mechanisms to make the optimization more stable. Finally, we see a potential to scaling up our method to a high number of particles to parallelize ES in an informed way. An investigation of scaling laws would be an intriguing avenue of research.
Acknowledgements.
This research was funded by the Amazon Development Center Germany GmbH. The authors want to thank anonymous reviewers for constructive feedback which helped improve the paper.
Supplementary Material The supplementary material is structured as follows:
  • Appendix A lists the vanilla CMA-ES algorithm for comparison to our method and the computation of all hyperparameters that we used.
  • Appendix B presents the full experimental details of our work.
  • Appendix C presents additional experimental results. These include ablations of SV-CMA-ES, an empirical runtime analysis, and an empirical convergence analysis.
Algorithm 2 The Vanilla CMA-ES Update. Adapted from Hansen [2016].
Input: Generation index $t$, search distribution parameters $\mathbf{x},\sigma,\mathbf{C}$; pop. size $n$, num. elites $m$; learning rates $\epsilon,\alpha_{\sigma},\alpha_{1},\alpha_{m},\alpha_{c}$; damping hyperparam. $d_{\sigma}$
1:Sample and evaluate new population of search points, for $i=1,\dots,n$
$\mathbf{y}_{i}=(\boldsymbol{\xi}_{i}-\mathbf{x})/\sigma,\quad\text{where }\boldsymbol{\xi}_{i}\sim{\cal N}(\mathbf{x},\sigma^{2}\mathbf{C})$
$f_{i}=f(\boldsymbol{\xi}_{i})$